# Long-term stable nonlinear evolutions of ultracompact black-hole mimickers

### v1

@@ Appendix C Radial Oscillation Frequencies
To obtain the first pulsation equation, we solve Eq. (15) for $\delta\dot{\psi}_{2}$ and substitute into Eq. (14) to obtain an equation that depends on $\delta\nu$ only through its radial derivative which, in turn, can be removed using (17). For the second pulsation equation, we combine Eqs. (16), (15) and (11) to obtain an equation that depends on $\delta\nu^{\prime}$ and $\delta\nu^{\prime\prime}$ but not $\delta\nu$ directly. By using Eq. (17), its derivative and Eq. (11), we can eliminate $\delta\nu$ completely and thus arrive at our final system of two ODEs,
@@
$f^{\prime\prime}=\,$ $X_{0}^{2}\bigg{[}2+\frac{2(A_{0}^{\prime})^{2}}{X_{0}^{2}A_{0}^{2}}+\frac{2\omega^{2}-\chi^{2}}{\alpha_{0}^{2}}+8\pi rA_{0}A_{0}^{\prime}-\frac{32A_{0}^{2}}{\sigma_{0}^{2}}\left(1+2\pi rA_{0}^{2}A_{0}^{\prime}\right)+\frac{8A_{0}^{4}}{\sigma_{0}^{4}}\left(9+12\pi rA_{0}A_{0}^{\prime}\right)\bigg{]}f$ (18)
$+\bigg{[}X_{0}^{2}\left(1-\frac{8A_{0}^{2}}{\sigma_{0}^{2}}+\frac{12A_{0}^{4}}{\sigma_{0}^{4}}\right)-\frac{1}{4\pi r^{2}A_{0}^{2}}-\frac{(A_{0}^{\prime})^{2}}{A_{0}^{2}}-\frac{A_{0}^{\prime}}{rA_{0}}-\frac{A_{0}^{\prime}\alpha_{0}^{\prime}}{A_{0}\alpha_{0}}+\frac{X_{0}^{\prime}}{X_{0}A_{0}}\left(A_{0}^{\prime}+\frac{1}{2\pi rA_{0}}\right)\bigg{]}g$
$+\left(\frac{X_{0}^{\prime}}{X_{0}}-\frac{\alpha_{0}^{\prime}}{\alpha_{0}}-\frac{2}{r}\right)f^{\prime}-\frac{1}{4\pi rA_{0}^{2}}g^{\prime},$
$g^{\prime\prime}=\,$ $32\pi rX_{0}^{2}\bigg{[}A_{0}A_{0}^{\prime}+\frac{\alpha_{0}^{\prime}A_{0}^{2}}{\alpha_{0}}-\frac{8A_{0}^{3}}{\sigma_{0}^{2}}\left(2A_{0}^{\prime}+\frac{A_{0}\alpha_{0}^{\prime}}{\alpha_{0}}\right)+\frac{X_{0}^{\prime}}{X_{0}}\left(A_{0}^{2}-\frac{4A_{0}^{4}}{\sigma_{0}^{2}}+\frac{6A_{0}^{6}}{\sigma_{0}^{4}}-\frac{(A_{0}^{\prime})^{2}}{X_{0}^{2}}\right)$ (19)
$+\frac{12A_{0}^{5}}{\sigma_{0}^{4}}\left(3A_{0}^{\prime}+\frac{A_{0}\alpha_{0}^{\prime}}{\alpha_{0}}\right)\bigg{]}f+3\left(\frac{X_{0}^{\prime}}{X_{0}}-\frac{\alpha_{0}^{\prime}}{\alpha_{0}}\right)g^{\prime}+\bigg{[}16\pi(A_{0}^{\prime})^{2}-\frac{\chi^{2}X_{0}^{2}-2(\alpha_{0}^{\prime})^{2}}{\alpha_{0}^{2}}-\frac{4\alpha^{\prime}_{0}}{r\alpha_{0}}+\frac{2}{r^{2}}$
$+\frac{2}{X_{0}}\left(X_{0}^{\prime\prime}+\frac{2\alpha_{0}^{\prime}X_{0}^{\prime}}{\alpha_{0}}-\frac{2(X_{0}^{\prime})^{2}}{X_{0}}-\frac{X_{0}^{\prime}}{r}\right)\bigg{]}g+16\pi rX_{0}^{2}A_{0}^{2}\left(1-\frac{8A_{0}^{2}}{\sigma_{0}^{2}}+\frac{12A_{0}^{4}}{\sigma_{0}^{4}}-2\frac{A_{0}A_{0}^{\prime}}{rX_{0}^{2}}\right)f^{\prime}.$
@@
Before we can solve this system numerically, we must determine its behavior in the limit $r\rightarrow 0$. Bearing in mind the even parity of scalar functions across $r=0$ and using our freedom to rescale the perturbations, we can expand them without loss of generality as $f=1+\frac{a}{2}r^{2}+\mathcal{O}(r^{4})$ and $g=\beta+\frac{\gamma}{2}r^{2}+\mathcal{O}(r^{4})$. Inserting into Eqs. (18) and (19) gives us
$a=$ $\frac{1}{3}\left(2+\frac{2\omega^{2}-\chi^{2}}{\alpha_{0}(0)^{2}}-\frac{32A_{0}(0)^{2}}{\sigma_{0}^{2}}+\frac{72A_{0}(0)^{4}}{\sigma_{0}^{4}}\right)$
$-\frac{\gamma}{8\pi A_{0}(0)^{2}},$ (20)
@@
as well as $\beta=0$ at leading order. There is a degeneracy, however, at the following order so that $\gamma$ remains undetermined. In consequence, we have two undetermined parameters, $\chi$ and $\gamma$, which we solve for using a shooting method.
@@
[…]
Combining with Eq. (15) allows us to eliminate $\delta\nu$ whence, using the background equations (11)-(13), we can write the integrand as a total derivative and find
@@
$\delta\mathcal{N}=\frac{1}{\omega X_{0}}\left(8\pi r^{2}\alpha_{0}A_{0}A_{0}^{\prime}f-r\alpha_{0}g\right)\Big{\rvert}^{\infty}_{r=0}.$ (22)
@@
Despite appearances, the first term is nonnegligible in this expression due to the exponential behavior of $f$ which vanishes only for precise choices of the shooting parameters $\chi$ and $\gamma$. Thus, for regular perturbations which remain finite at $r=0$, the condition $\delta N=0$ supplies a second boundary condition at infinity.
Finally, it is worth noting that because $\delta\lambda$ is related to the perturbation to the total mass according to
$\delta\lambda=\frac{2X_{0}^{2}}{r}\delta M,$ (23)
@@
at large radius, our two conditions amount to enforcing that both the mass and particle number of the BS are unchanged by the perturbation. In practice, we solve Eqs. (18) and (19) using a standard fourth-order Runge-Kutta method, shooting for the undetermined values $\chi$ and $\gamma$ for each model. We determine these two parameters using a bisection method similar to that commonly used to determine the frequency $\omega$ for equilibrium BS models. Interestingly, we encounter the same challenging accuracy requirements in our determination of $\chi$ and $\gamma$ for thin-shell BSs as observed when calculating the BS frequency $\omega$; therefore, we also compute the radial oscillation spectrum using quadruple precision.
@@
In this work, we restrict our attention to presenting positive $\chi_{0}^{2}$ values corresponding to radially stable models, though this approach can be used to compute negative values as well as higher-order oscillation frequencies. The point in the BS parameter space at which models transition from stable to unstable is marked by the existence of a zero-frequency perturbation, $\chi_{0}^{2}=0$. Such zero-frequency perturbations can only exist where there are two neighboring solutions with the same mass (or equivalently, particle number). Therefore, the condition $\mathrm{d}M/\mathrm{d}A(0)=0$ is a necessary but insufficient criterion for the transition of a radial oscillation mode from stability to instability. For mini BSs, only models with smaller central amplitude than that corresponding to the maximum mass model are stable Gleiser and Watkins (1989). There is hence only a single stable branch for mini BSs. However, for solitonic models with sufficiently small $\sigma_{0}$, a second stable branch emerges; see e.g. Fig. 4 in Ref. Ge et al. (2024) for representative examples.
In Fig. 12, we present values of $\chi_{0}^{2}$ for mini and solitonic BS models on both stable branches. Our results confirm the radial stability of solitonic BS models on the second stable branch, including in the thin-shell regime where $M$ and $\omega$ become multi-valued functions of the central amplitude
Indeed, we see that the curve of $\chi_{0}^{2}$ against the central amplitude also becomes multivalued for these $\sigma_{0}$-values, and in fact self-intersects. In the case of mini BSs, our results are compatible with those computed in Ref. Kain (2021), while the solitonic results are verified by comparing to the power spectrums of dynamical evolutions.
@@
Figure 12: Fundamental radial oscillation frequencies $\chi_{0}^{2}$ against central amplitude $A_{0}$ for mini and solitonic BSs with a variety of $\sigma_{0}$-values. The main figure shows the second stable branch, while the inset shows the first. Central amplitudes are rescaled by the appropriate $\sigma_{0}$-value, apart from mini BSs, which undergo no rescaling.
Appendix D Adiabatic Effective Potential
@@
To assess whether the BS remains ultracompact throughout the evolution, we need to check whether light rings persist. As the BS is evolved in 3+1 dimensions, the spacetime generally loses the spherical symmetry as a result of numerical errors and gauge effects that excite asymmetric modes. In terms of the 3+1 metric
$\mathrm{d}s^{2}=-\alpha^{2}\mathrm{d}t^{2}+\gamma_{ij}\left(\mathrm{d}x^{i}+\beta^{i}\mathrm{d}t\right)\left(\mathrm{d}x^{j}+\beta^{j}\mathrm{d}t\right)\,,$ (24)
@@
$H(r,\theta)=\sqrt{\frac{-g_{tt}}{g_{\phi\phi}}}\,$ (25)
is only defined for stationary spacetimes that exhibit spherical symmetry, whose metric components we henceforth denote $g_{\alpha\beta}$ in this section. Given that our evolutions are expected to approximately respect the symmetries of the initial data, the presence of (stable) bound geodesics can still be deduced from the Adiabatic Effective Potential (AEP).
@@
We adapt the approach of Ref. Cunha et al. (2023) from (approximate) axisymmetry to our spherically symmetric setting. Under the assumptions outlined in this work, the metric component $g_{tt}$ in the effective potential (25) can be replaced by the evolution variables according to
$g_{tt}~{}~{}\rightarrow~{}~{}-\langle\alpha^{2}\rangle+\langle\gamma_{ij}\beta^{i}\beta^{j}\rangle\,,$ (26)
@@
where the brackets denote averaging over a sphere of fixed coordinate radius $R=\sqrt{x^{2}+y^{2}+z^{2}}$ and $x,y,z$ are the Cartesian coordinates of the evolution grid. We convert $R$ into the areal radius through [^1] Note that it is not guaranteed that the coordinate sphere actually corresponds to a sphere in terms of the areal radius. Again, if the evolution remains sufficiently close to spherical symmetry, this is a good approximation.
$r^{2}=\frac{\mathcal{A}}{4\pi}=\frac{1}{4\pi}\langle\sqrt{\text{det }}q\rangle$ (27)
@@
where $q$ is the metric induced on the sphere $R=\mathrm{const}$. Mapping the Cartesian evolution coordinates $(x,y,z)$ to angular coordinates $(R,\tilde{\theta},\tilde{\phi})$ according to
$x=R\sin\tilde{\theta}\cos\tilde{\phi}\,,~{}~{}\\ y=R\sin\tilde{\theta}\sin\tilde{\phi}\,,~{}~{}\\ z=R\cos\tilde{\theta}\,,$
we can express the induced metric as
$q_{\tilde{\theta}\tilde{\theta}}=\,$ $z^{2}(\gamma_{xx}\cos^{2}\tilde{\phi}+\gamma_{yy}\sin^{2}\tilde{\phi})+\gamma_{zz}R^{2}\sin^{2}\tilde{\theta}$
$+\gamma_{xy}z^{2}\sin(2\tilde{\phi})-2zR\sin\tilde{\theta}(\gamma_{xz}\cos\tilde{\phi}+\gamma_{yz}\sin\tilde{\phi})\,,$
$q_{\tilde{\phi}\tilde{\phi}}=\,$ $\gamma_{xx}y^{2}-2\gamma_{xy}xy+\gamma_{yy}x^{2}\,,$
$q_{\tilde{\theta}\tilde{\phi}}=\,$ $\gamma_{yy}xz\sin\tilde{\phi}-\gamma_{xx}yz\cos\tilde{\phi}+\gamma_{xy}z(x\cos\tilde{\phi}-y\sin\tilde{\phi})$
$+R\sin\tilde{\theta}(\gamma_{xz}y-\gamma_{yz}x)\,.$
@@
In the equatorial plane $\theta=\frac{\pi}{2}$ of the stationary and spherically symmetric spacetime, we have $g_{\phi\phi}=r^{2}$. Replacing $g_{\phi\phi}\rightarrow\frac{1}{4\pi}\langle\sqrt{\det q}\rangle$ and using Eqs. (25)-(27) we now define the AEP as
@@
$H_{\text{eff}}(R)=\sqrt{\frac{\langle\alpha^{2}\rangle-\langle\gamma_{ij}\beta^{i}\beta^{j}\rangle}{\langle\sqrt{\text{det }q}\rangle/4\pi}}\,,$ (28)
@@
which is straightforwardly translated into a function of areal radius using Eq. (27). In the idealized case where our simulation is perfectly spherically symmetric and stationary, Eq. (28) reduces to Eq. (25). As long as this AEP has extrema we have evidence that the BS is still ultracompact, featuring a pair of (perturbed) light rings. In practice, the AEP is extracted at a number of coordinate radii with resolution $\Delta R\approx 0.1$ in the range where we expect the light rings to be.
@@
$\delta(X)\equiv\frac{\sigma(X)}{\langle X\rangle}\equiv\frac{\sqrt{\langle\left(X-\langle X\rangle\right)^{2}\rangle}}{\langle X\rangle}\,,$ (29)
@@
where $\langle X\rangle$ denotes the average over the coordinate sphere with radius $R$. Figure 13 shows the relative standard deviation for the individual terms of Eq. (28), as well as that of the AEP itself, as obtained for 192 points on each sphere.
In the case of $\beta^{2}\equiv\gamma_{ij}\beta^{i}\beta^{j}$ we have $\delta(\beta^{2})\gtrsim 10^{-2}$, reaching order unity at certain times. This relatively large $\delta(\beta^{2})$ is merely due to the shift vector having a very small norm. This is further underlined by the ratio $\langle\beta^{2}\rangle/\langle\alpha^{2}\rangle\lesssim 10^{-9}$. For the other variables, the relative standard deviation is $\lesssim 10^{-6}$, suggesting that the spacetime is very close to spherical symmetry and approximately stationary.
In order to investigate whether these variations influence the conclusion about the presence of light rings, we plot $\sigma(H_{\text{eff}})/\Delta H$, where $\Delta H$ is the potential difference between the extracted extrema of the AEP. This value varies around $\sim 10^{-3}-10^{-2}$, implying that the deviations from our symmetry assumptions explain at most 1% of the observed potential difference between the extrema, demonstrating their robustness.
@@
Figure 13: Diagnostics to assess the deviations from spherical symmetry and stationarity for BS model S08A06. Brackets denote averaging over an extraction sphere, whereas $\sigma$ denotes the standard deviation with respect to this average value. The relative standard deviation is $\delta(X)\equiv\sigma(X)/\langle X\rangle$, and $\Delta H$ denotes the potential difference between the two extrema of the AEP. The lines with markers correspond to time $\mu t=75$, after the inital dynamics start to settle down. The other lines correspond to extractions at later times ($\mu t\in\{250,500,\dots,2000\}$) to illustrate that the conclusions also hold at later times.
@@
The conclusions for S06A044 are similar and displayed in Fig. 14. The major difference is that the norm of the shift vector compared to the lapse is comparatively larger at $\langle\beta^{2}\rangle/\langle\alpha^{2}\rangle\sim 10^{-6}$. This is due to the larger dynamics in this more challenging simulation, as was apparent from other diagnostics. The ratio $\sigma(H_{\text{eff}})/\Delta H\lesssim 10^{-3}$ is smaller by an order of magnitude, an artifact of the more pronounced extrema for this more extreme model.
Figure 14: Same as Fig. 13, but for BS model S06A044.
Appendix E Ray tracing

### v2

@@ III Radial Oscillation Frequencies
To obtain the first pulsation equation, we solve Eq. (S9) for $\delta\dot{\psi}_{2}$ and substitute into Eq. (S8) to obtain an equation that depends on $\delta\nu$ only through its radial derivative which, in turn, can be removed using Eq. (S11). For the second pulsation equation, we combine Eqs. (S6), (S9) and (S10) to obtain an equation that depends on $\delta\nu^{\prime}$ and $\delta\nu^{\prime\prime}$ but not $\delta\nu$ directly. By using Eq. (S11), its derivative and Eq. (S6), we can eliminate $\delta\nu$ completely and thus arrive at our final system of two ODEs,
@@
$f^{\prime\prime}=\,$ $X_{b}^{2}\bigg{[}2+\frac{2(A_{b}^{\prime})^{2}}{X_{b}^{2}A_{b}^{2}}+\frac{2\omega^{2}-\chi^{2}}{\alpha_{b}^{2}}+8\pi rA_{b}A_{b}^{\prime}-\frac{32A_{b}^{2}}{\sigma_{0}^{2}}\left(1+2\pi rA_{b}^{2}A_{b}^{\prime}\right)+\frac{8A_{b}^{4}}{\sigma_{0}^{4}}\left(9+12\pi rA_{b}A_{b}^{\prime}\right)\bigg{]}f$ (S12)
$+\bigg{[}X_{b}^{2}\left(1-\frac{8A_{b}^{2}}{\sigma_{0}^{2}}+\frac{12A_{b}^{4}}{\sigma_{0}^{4}}\right)-\frac{1}{4\pi r^{2}A_{b}^{2}}-\frac{(A_{b}^{\prime})^{2}}{A_{b}^{2}}-\frac{A_{b}^{\prime}}{rA_{b}}-\frac{A_{b}^{\prime}\alpha_{b}^{\prime}}{A_{b}\alpha_{b}}+\frac{X_{b}^{\prime}}{X_{b}A_{b}}\left(A_{b}^{\prime}+\frac{1}{2\pi rA_{b}}\right)\bigg{]}g$
$+\left(\frac{X_{b}^{\prime}}{X_{b}}-\frac{\alpha_{b}^{\prime}}{\alpha_{b}}-\frac{2}{r}\right)f^{\prime}-\frac{1}{4\pi rA_{b}^{2}}g^{\prime},$
$g^{\prime\prime}=\,$ $32\pi rX_{b}^{2}\bigg{[}A_{b}A_{b}^{\prime}+\frac{\alpha_{b}^{\prime}A_{b}^{2}}{\alpha_{b}}-\frac{8A_{b}^{3}}{\sigma_{0}^{2}}\left(2A_{b}^{\prime}+\frac{A_{b}\alpha_{b}^{\prime}}{\alpha_{b}}\right)+\frac{X_{b}^{\prime}}{X_{b}}\left(A_{b}^{2}-\frac{4A_{b}^{4}}{\sigma_{0}^{2}}+\frac{6A_{b}^{6}}{\sigma_{0}^{4}}-\frac{(A_{b}^{\prime})^{2}}{X_{b}^{2}}\right)$ (S13)
$+\frac{12A_{b}^{5}}{\sigma_{0}^{4}}\left(3A_{b}^{\prime}+\frac{A_{b}\alpha_{b}^{\prime}}{\alpha_{b}}\right)\bigg{]}f+3\left(\frac{X_{b}^{\prime}}{X_{b}}-\frac{\alpha_{b}^{\prime}}{\alpha_{b}}\right)g^{\prime}+\bigg{[}16\pi(A_{b}^{\prime})^{2}-\frac{\chi^{2}X_{b}^{2}-2(\alpha_{b}^{\prime})^{2}}{\alpha_{b}^{2}}-\frac{4\alpha^{\prime}_{0}}{r\alpha_{b}}+\frac{2}{r^{2}}$
$+\frac{2}{X_{b}}\left(X_{b}^{\prime\prime}+\frac{2\alpha_{b}^{\prime}X_{b}^{\prime}}{\alpha_{b}}-\frac{2(X_{b}^{\prime})^{2}}{X_{b}}-\frac{X_{b}^{\prime}}{r}\right)\bigg{]}g+16\pi rX_{b}^{2}A_{b}^{2}\left(1-\frac{8A_{b}^{2}}{\sigma_{0}^{2}}+\frac{12A_{b}^{4}}{\sigma_{0}^{4}}-2\frac{A_{b}A_{b}^{\prime}}{rX_{b}^{2}}\right)f^{\prime}.$
@@
Before we can solve this system numerically, we must determine its behavior in the limit $r\rightarrow 0$. Bearing in mind the even parity of scalar functions across $r=0$ and using our freedom to rescale the perturbations, we can expand them without loss of generality as $f=1+\frac{a}{2}r^{2}+\mathcal{O}(r^{4})$ and $g=\beta+\frac{\gamma}{2}r^{2}+\mathcal{O}(r^{4})$. Inserting into Eqs. (S12) and (S13) gives us
$a=$ $\frac{1}{3}\left(2+\frac{2\omega^{2}-\chi^{2}}{\alpha_{b}(0)^{2}}-\frac{32A_{b}(0)^{2}}{\sigma_{0}^{2}}+\frac{72A_{b}(0)^{4}}{\sigma_{0}^{4}}\right)$
$-\frac{\gamma}{8\pi A_{b}(0)^{2}},$ (S14)
@@
as well as $\beta=0$ at leading order. There is a degeneracy, however, at the following order so that $\gamma$ remains undetermined. In consequence, we have two undetermined parameters, $\chi$ and $\gamma$, which need to be finetuned to obtain regular spacetimes.
@@
$\delta\mathcal{N}=4\pi\int_{0}^{\infty}\frac{A_{b}^{2}r^{2}}{\alpha_{b}}dr\left[\omega X_{b}(\delta\lambda-\delta\nu)+2\delta\dot{\psi}_{2}+4\omega\delta\psi_{1}\right]\,.$
Combining with Eq. (S9) allows us to eliminate $\delta\nu$ whence, using the background equations (S2)-(S6), we can write the integrand as a total derivative and find
$\delta\mathcal{N}=\frac{1}{\omega X_{b}}\left(8\pi r^{2}\alpha_{b}A_{b}A_{b}^{\prime}f-r\alpha_{b}g\right)\Big{\rvert}^{\infty}_{r=0}.$ (S15)
@@
Despite appearances, the first term is nonnegligible in this expression due to the exponential behavior of $f$ which vanishes only for precise choices of the shooting parameters $\chi$ and $\gamma$. Thus, for regular perturbations which remain finite at $r=0$, the condition $\delta N=0$ supplies a second boundary condition at infinity. Finally, it is worth noting that because $\delta\lambda$ is related to the perturbation of the total mass,
$\delta\lambda=\frac{2X_{b}^{2}}{r}\delta M,$ (S16)
@@
at large radius, our two conditions amount to enforcing that both the mass and particle number of the BS are unchanged by the perturbation. In practice, we solve Eqs. (S12) and (S13) using a standard fourth-order Runge-Kutta method, shooting for the undetermined values $\chi$ and $\gamma$ for each model. We determine these two parameters using a bisection method similar to that commonly used to determine the frequency $\omega$ for equilibrium BS models. Interestingly, we encounter the same challenging accuracy requirements in our determination of $\chi$ and $\gamma$ for thin-shell BSs as observed when calculating the BS frequency $\omega$; therefore we also compute the radial oscillation spectrum using quadruple precision.
@@
With this method, we can compute real and imaginary $\chi_{0}$ as well as overtone frequencies. In this work, we wish to identify stable systems and therefore restrict our attention to identifying regimes of positive versus negative $\chi_{0}^{2}$. The point in the BS parameter space at which models transition from stable to unstable is marked by the existence of a zero-frequency perturbation, $\chi_{0}^{2}=0$. Such zero-frequency perturbations can only exist where two neighboring solutions with the same mass (or, equivalently, Noether charge) exist. The condition $\mathrm{d}M/\mathrm{d}A_{0}=0$ is therefore a necessary but insufficient criterion for the transition of a radial oscillation mode from stability to instability. For mini BSs, there exists only one stable branch consisting of models with central amplitude smaller than that corresponding to the maximum-mass model Gleiser and Watkins (1989). However, for solitonic models with sufficiently small $\sigma_{0}$, a second stable branch emerges; see e.g. Fig. 4 in Ref. Ge et al. (2024) for representative examples.
We illustrate the existence of two stable branches for solitonic BSs in Fig. S6. There, for each family of BSs, a positive $\chi_{0}^{2}>0$ is obtained across two disjoint regions in the parameter space, including the ultracompact thin-shell models in the regime where $M$ and $\omega$ are multi-valued functions of $A_{0}$. Indeed, the curve $\chi^{2}_{0}(A_{0})$ also becomes multivalued for these $\sigma_{0}$-values, and self-intersects. In the case of mini BSs, our results are compatible with those computed in Ref. Kain (2021).
@@
Figure S6: Fundamental radial oscillation frequencies $\chi_{0}^{2}$ against central amplitude $A_{0}$ for mini and solitonic BSs with a variety of $\sigma_{0}$-values. The main figure shows the second stable branch, while the inset shows the first. Central amplitudes are rescaled by the appropriate $\sigma_{0}$-value, apart from mini BSs, which undergo no rescaling.
IV Adiabatic Effective Potential
@@
To assess whether our BSs remains ultracompact throughout their evolution, we need to check whether the light-ring structure persists. As the BS is evolved in 3+1 dimensions, the spacetime generally loses its spherical symmetry as a result of numerical errors and gauge effects that excite asymmetric modes. In terms of the 3+1 metric
$\mathrm{d}s^{2}=-\alpha^{2}\mathrm{d}t^{2}+\gamma_{ij}\left(\mathrm{d}x^{i}+\beta^{i}\mathrm{d}t\right)\left(\mathrm{d}x^{j}+\beta^{j}\mathrm{d}t\right)\,$, (S17)
@@
$H(\tilde{r},\tilde{\theta})=\sqrt{\frac{-g_{\tilde{t}\tilde{t}}}{g_{\tilde{\phi}\tilde{\phi}}}}\,$ (S18)
is only defined for stationary spacetimes that exhibit spherical symmetry; here a tilde distinguishes coordinates of such a spacetime from those used in our approximately symmetric numerical evolutions.
@@
Given that our evolutions are expected to approximately respect the symmetries of the initial data, the presence of (stable) bound geodesics can still be deduced from an Adiabatic Effective Potential (AEP). To derive the AEP, we adapt the approach of Ref. Cunha et al. (2023) from (approximate) axisymmetry and apply it to our spherically symmetric setting. Under the assumptions outlined in that work, the metric component $g_{\tilde{t}\tilde{t}}$ in the effective potential (S18) can be replaced by the evolution variables of the 3+1 line element (S17) according to
$g_{\tilde{t}\tilde{t}}~~\rightarrow~~-\langle\alpha^{2}\rangle+\langle\gamma_{ij}\beta^{i}\beta^{j}\rangle.$ (S19)
@@
Here the brackets denote averaging over a sphere of fixed coordinate radius $R=\sqrt{x^{2}+y^{2}+z^{2}}$ and $x,y,z$ are the Cartesian coordinates of the evolution grid. We convert $R$ into the areal radius $r$ by calculating the proper area $\mathcal{A}$ of the sphere [^1] Note that it is not guaranteed that the coordinate sphere actually corresponds to a sphere in terms of the areal radius. Again, if the evolution remains sufficiently close to spherical symmetry, this is a good approximation.
$r^{2}=\frac{\mathcal{A}}{4\pi}=\frac{1}{4\pi}\oint_{S^{2}}\sqrt{\det q}\>\mathrm{d}\theta\>\mathrm{d}\phi\,$, (S20)
@@
where $q$ is the metric induced on the sphere. Mapping between Cartesian coordinates $(x,y,z)$ and angular coordinates $(R,\theta,\phi)$, the induced metric becomes
$q_{\theta\theta}=\,$ $z^{2}(\gamma_{xx}\cos^{2}\phi+\gamma_{yy}\sin^{2}\phi)+\gamma_{zz}R^{2}\sin^{2}\theta$
$+\gamma_{xy}z^{2}\sin(2\phi)-2zR\sin\theta(\gamma_{xz}\cos\phi+\gamma_{yz}\sin\phi)\,$,
$q_{\phi\phi}=\,$ $\gamma_{xx}y^{2}-2\gamma_{xy}xy+\gamma_{yy}x^{2}\,$,
$q_{\theta\phi}=\,$ $\gamma_{yy}xz\sin\phi-\gamma_{xx}yz\cos\phi+\gamma_{xy}z(x\cos\phi-y\sin\phi)$
$+R\sin\theta(\gamma_{xz}y-\gamma_{yz}x)\,.$
@@
In the equatorial plane $\tilde{\theta}=\frac{\pi}{2}$ of the stationary and spherically symmetric spacetime, we have $g_{\tilde{\phi}\tilde{\phi}}=\tilde{r}^{2}$. Replacing $\tilde{r}\rightarrow r$ and using Eqs. (S18)-(S20) we now define the AEP as
@@
$H_{\text{eff}}(r)=\sqrt{\frac{\langle\alpha^{2}\rangle-\langle\gamma_{ij}\beta^{i}\beta^{j}\rangle}{r}}\,.$ (S21)
@@
In the idealized case where our simulation is perfectly spherically symmetric and stationary, Eq. (S21) reduces to Eq. (S18). As long as this AEP has extrema we have evidence that the BS is still ultracompact, featuring a pair of (perturbed) light rings. In practice, we extract the AEP at a number of coordinate radii with resolution $\Delta R\approx 0.1$ in the range where we expect the light rings to be.
@@
$\delta(X)\equiv\frac{\sigma(X)}{\langle X\rangle}\equiv\frac{\sqrt{\langle\left(X-\langle X\rangle\right)^{2}\rangle}}{\langle X\rangle}\,$, (S22)
@@
Figure S7 shows the relative standard deviation for the individual terms of Eq. (S21), as well as that of the AEP itself, as obtained for 192 points on each sphere.
The largest relative deviation from spherical symmetry is seen for $\beta^{2}\equiv\gamma_{ij}\beta^{i}\beta^{j}$, where $\delta(\beta^{2})\gtrsim 10^{-2}$, reaching order unity at some times; this is merely a consequence of our definition (S22) and the near-zero average $\langle\beta^{2}\rangle$. This interpretation is supported by the very small ratio $\langle\beta^{2}\rangle/\langle\alpha^{2}\rangle\lesssim 10^{-9}$. The relative standard deviation of the remaining variables is $\lesssim 10^{-6}$, suggesting that the spacetime is very close to spherical symmetry and approximately stationary. We further investigate whether these small variations have any influence on the light rings’ presence and structure. For this purpose we also display $\sigma(H_{\text{eff}})/\Delta H$ as the purple curve in Fig. S7, where $\Delta H$ is the potential difference between the extracted extrema of the AEP. Its value varies around $\sim 10^{-3}-10^{-2}$, implying that the deviations from our symmetry assumptions explain at most 1% of the observed potential difference between the extrema.
@@
Figure S7: Diagnostics to assess the deviations from spherical symmetry and stationarity for S08A06. Brackets denote averaging over an extraction sphere, whereas $\sigma$ denotes the standard deviation from this average value. The relative standard deviation is $\delta(X)\equiv\sigma(X)/\langle X\rangle$, and $\Delta H$ denotes the potential difference between the two extrema of the AEP. The lines with markers correspond to time $\mu t=75$, after early gauge dynamics have settled down. The other lines correspond to extractions at later times ($\mu t\in\{250,500,\dots,2000\}$) to illustrate that the conclusions also hold at later times.
@@
The conclusions for S06A044 are similar and displayed in Fig. S8. The major difference is that $\langle\beta^{2}\rangle/\langle\alpha^{2}\rangle\sim 10^{-6}$ is considerably larger now. This is in line with the more pronounced oscillations also observed in other diagnostics for this numerically more challenging configuration; cf. Figs. 2 and 4 in the main text. The ratio $\sigma(H_{\text{eff}})/\Delta H\lesssim 10^{-3}$, however, is smaller by an order of magnitude, a side effect of the more pronounced extrema in the potential for this more extreme model.
Figure S8: Same as Fig. S7, but for BS model S06A044.
V Ray tracing